# Can We Predict Before Executing Machine Learning Agents?

### v1

@@ C.1 Fine-grained Performance on Prediction Corpus
C.2 Detailed Performance Metrics of ForeAgent on AI4Science Benchmarks
@@
C.3 Search Efficiency Analysis of ForeAgent
@@
C.4 Decision Fidelity Analysis of ForeAgent
To audit decision quality without accessible ground truth for skipped nodes (which yield no test metrics), we evaluate the trajectory consistency of executed steps by aligning internal Validation Scores ($S_{val}$) with external Test Scores ($S_{test}$). We define two metrics: Solution Evolution Consistency, which checks if the validation gains between consecutive executed steps ($S_{val}^{B}>S_{val}^{A}$) are consistent with test outcomes, and Global Pairwise Consistency, which measures the ranking agreement across the entire search history.
As shown in Table 13, ForeAgent exhibits a marginal decrease in Solution Evolution Consistency compared to the execution-based baseline AIDE ($0.756$ vs. $0.779$). This slight trade-off is acceptable given the substantial efficiency gains. Conversely, ForeAgent achieves superior Global Pairwise Consistency ($0.801$ vs. $0.741$), suggesting that our implicit evaluation acts as a regularizer, effectively filtering out overfitting candidates to yield a more stable search trajectory.
C.5 Licensing and Artifact Usage
@@
C.6 Computational Infrastructure and Budget
@@
C.7 Software Dependencies and Metric Implementation
@@
Task / Aggregation Metric: Average Accuracy Solution Evolution Consistency Global Pairwise Consistency
AIDE (Baseline) ForeAgent AIDE (Baseline) ForeAgent
Stanford Covid Vaccine $\bm{0.950}_{\pm\bm{0.030}}$ $0.644_{\pm 0.258}$ $\bm{0.817}_{\pm\bm{0.067}}$ $0.679_{\pm 0.179}$
Statoil Iceberg Classifier $\bm{0.975}_{\pm\bm{0.043}}$ $0.840_{\pm 0.073}$ $0.792_{\pm 0.036}$ $\bm{0.813}_{\pm\bm{0.024}}$
Ventilator Pressure Prediction $0.750_{\pm 0.000}$ $\bm{0.905}_{\pm\bm{0.165}}$ $0.393_{\pm 0.556}$ $\bm{0.926}_{\pm\bm{0.128}}$
Aerial Cactus Identification* $0.348_{\pm 0.399}$ $\bm{0.472}_{\pm\bm{0.411}}$ $\bm{0.758}_{\pm\bm{0.108}}$ $0.696_{\pm 0.125}$
Histopathologic Cancer Detection* $\bm{1.000}_{\pm\bm{0.000}}$ $\bm{1.000}_{\pm\bm{0.000}}$ $\bm{1.000}_{\pm\bm{0.000}}$ $0.889_{\pm 0.193}$
Overall Average $\bm{0.779}_{\pm\bm{0.337}}$ $0.756_{\pm 0.280}$ $0.741_{\pm 0.249}$ $\bm{0.801}_{\pm\bm{0.159}}$
Table 13: Decision Fidelity Analysis (AIDE vs. ForeAgent). We report the mean and standard deviation using the format mean (std). Solution Evolution Consistency measures the reliability of decisions along the iterative evolution chain, while Global Pairwise Consistency measures the ranking quality of the entire search trajectory. The best results are highlighted in bold.
@@ Appendix D Detailed Qualitative Analysis
To validate the model’s reasoning depth and provide transparency into our pipeline, we present three qualitative examples. First, we analyze a reasoning trajectory in the Google Quest Challenge to illustrate how the model overcomes human bias (Finding 5). Subsequently, we provide visual samples of two critical system artifacts: the Verbal Data Report and the Task Instruction, enabling a concrete inspection of the agent’s input and context.
@@ D.1 Case I: Overcoming Complexity Bias (Reasoning Analysis)
D.2 Case II: Sample of the Verbal Data Report
@@
D.3 Case III: Sample of the Task Instruction ($I$)

### v2

@@ C.1 Fine-grained Performance on Prediction Corpus
C.2 Analysis of Pair Source and Trajectory Variance
To rigorously test whether random sampling introduces easy pairs by comparing an initial half baked script versus a refined solution from the same trajectory, we performed a stratified analysis. We split our test set into two subsets: Within Trajectory Pairs (both solutions come from the same agent run session) and Cross Trajectory Pairs (solutions come from different agent run sessions or different tasks).
As shown in table 11, our experimental results indicate that performance inflation is not observed between two subsets. For DeepSeek V3.2, accuracy on Within Trajectory pairs (60.4%) is slightly lower than on Cross Trajectory pairs (61.7%). For GPT 5.1, the performance is statistically identical across both subsets.
A qualitative inspection of the trajectories reveals why the assumption of early half baked code does not hold for our specific agents (AIDE and AutoMind). Unlike human developers who may write broken snippets in early stages, these autonomous agents are prompted to generate fully executable end to end scripts at every step of the iteration. Even the first generated solution is typically a complete runnable pipeline. Later steps represent methodological refinements, such as switching algorithms or adding feature engineering, rather than fixing broken code.
Therefore, distinguishing between an early complete script and a late complete script relies on subtle algorithmic reasoning, not on detecting obvious syntax errors or incompleteness. The model’s predictive capability is robust and reflects genuine reasoning about solution quality rather than the exploitation of trajectory artifacts.
C.3 Detailed Performance Metrics of ForeAgent on AI4Science Benchmarks
@@
C.4 Search Efficiency Analysis of ForeAgent
@@
C.5 Decision Fidelity and Reliability in Local Iterations
To investigate whether the framework maintains predictive reliability on noisy intermediate code and prevents the agent from going deeper into incorrect trajectories, we conducted a granular analysis of the decision making process comparing AIDE (execution only) and ForeAgent across 15 runs. We focused on three key metrics. Test Improve Rate is the probability that a modification intended to improve the code based on Validation feedback actually yields a better score on the hidden Test set. Test Non degrade Rate is the probability that the test score of the new candidate is greater than or equal to its parent, indicating the modification did not harm the performance. Val vs Test Agreement measures how often the Validation signal correctly predicts the Test direction for iterative improvement steps.
As shown in table 14, the results directly address the impact of intermediate noise. The execution only baseline (AIDE) struggles with non finalized code, showing a low success rate of 30.39%. This indicates that in local iterations, validation scores are extremely noisy. Traditional agents frequently go deeper into bad trajectories because they are misled by validation overfitting.
ForeAgent significantly improves the Test Improve Rate to 53.49%. By introducing the World Model as a filter before execution, the framework successfully prunes many candidate nodes that achieve high validation scores but are logically flawed. This proves that the World Model does not regress to the baseline when faced with draft code. Instead, it acts as a robust Semantic Safeguard, effectively filtering through the noise of intermediate development and reducing the risk of following incorrect trajectory paths.
Although the Val vs Test Agreement for local improve pairs shows a slight decrease (75.23%), it remains highly consistent with the global 72% theoretical ceiling discussed in Section 5.4. The higher actual success rate demonstrates that ForeAgent is a more reliable navigator for autonomous research and development than execution only feedback.
C.6 Ablation Study on Top $k$ Selection
We evaluated the parameter $k$ to balance search breadth and selection stability, comparing our default $k=1$ with $k=2$ and AIDE baseline.
Table 15 reveals three findings: (1) Search Space Expansion: Setting $k=2$ forces the execution of two candidates per iteration, surging the average node count to 157.87. (2) Metric Correlation: ForeAgent ($k=1$) balances stable Val Test Agreement (75.23%) with a substantially improved Test Improve Rate (53.49%). (3) Performance Degradation: Increasing $k$ amplifies exposure to noisy validation signals and increases variance, which destabilizes exploration and lowers the overall score.
In conclusion, $k=1$ provides the optimal balance. We note this preliminary study only isolates $k$. Future work will comprehensively explore tuning additional hyperparameter combinations.
C.7 Licensing and Artifact Usage
@@
C.8 Computational Infrastructure and Budget
@@
C.9 Software Dependencies and Metric Implementation
@@
Model Original (Full Set) Within Trajectory Cross Trajectory
DeepSeek V3.2 61.5% 60.4% 61.7%
GPT 5.1 58.8% 58.7% 58.9%
Table 11: Performance Breakdown by Pair Source. Accuracy on Within Trajectory pairs is not inflated.
@@
Metric AIDE (Execution only) ForeAgent (Ours)
Test Improve Rate 30.39% $\pm$ 11.93% 53.49% $\pm$ 14.70%
Test Non degrade Rate 30.50% $\pm$ 11.75% 53.97% $\pm$ 15.00%
Val vs Test Agreement (Local) 78.87% $\pm$ 10.19% 75.23% $\pm$ 13.01%
Table 14: Decision Reliability in Local Iterations. We report the mean and standard deviation across runs.
Model Mean Score Node Count Val Test Agreement Test Improve Rate
AIDE (No WM) 0.695 $\pm$ 0.298 44.00 $\pm$ 33.95 78.87% $\pm$ 10.19% 30.39% $\pm$ 11.93%
ForeAgent ($k=1$) 0.739 $\pm$ 0.295 99.60 $\pm$ 71.49 75.23% $\pm$ 13.01% 53.49% $\pm$ 14.70%
ForeAgent ($k=2$) 0.642 $\pm$ 0.353 157.87 $\pm$ 107.11 64.15% $\pm$ 26.57% 35.47% $\pm$ 23.00%
Table 15: Ablation Results on Top $k$ Selection
@@ Appendix D Detailed Qualitative Analysis
Case Study: Prioritizing Domain Fit over SOTA Architecture Task: TGS Salt Identification (Pixel-level Segmentation). Data Profile: Small scale ($101\times 101$) grayscale seismic images with weak intensity contrast and high variability in mask coverage. Data Report Insight: High sensitivity to spatial artifacts; resizing introduces interpolation noise that destroys local salt boundary details. Depth data is unavailable (NaN). Solution 0: Vision Transformer Stack: ViT-B/16 (Pre-trained on ImageNet) Setup: Resize to $224\times 224$, BCE + Dice Loss, attempts to use depth channel. Actual Score: 0.1301 (mAP)  Human Intuition "ViT is a top-tier SOTA model offering global context modeling. Pre-training on ImageNet guarantees strong feature extraction capabilities." Verdict: Strong Favorite World Model Critique "Resizing $101\times 101$ inputs to $224\times 224$ risks losing fine details. Domain gap between ImageNet (RGB) and seismic data (grayscale) is severe. Incorrectly relies on missing depth data." Verdict: Resolution Mismatch Solution 1: Standard U-Net Stack: Custom U-Net with Hybrid Attention Setup: Native $101\times 101$ resolution, BCE Loss, Test Time Augmentation (TTA). Actual Score: 0.6396 (mAP)  Human Intuition "Standard U-Net is an older, basic baseline. It lacks the global receptive field and advanced attention mechanisms of modern transformers." Verdict: Weak Baseline World Model Insight "Preserves native spatial resolution without interpolation noise. Attention mechanisms efficiently capture multi-scale features and spatial dependencies crucial for weak-contrast segmentation." Verdict: Optimal Fit   Outcome: Solution 1 outperformed Solution 0. The World Model was not swayed by the “SOTA” status, correctly prioritizing Native Spatial Resolution and Domain Fit over superficial architectural complexity.
Figure 11: Case Study: Domain Fit vs. Architectural Sophistication. This example highlights a “square peg in a round hole” scenario. While human intuition might favor the SOTA Vision Transformer (Solution 0), the World Model detects that forcing small $101\times 101$ seismic images into ViT’s required resolution introduces destructive interpolation noise. It rejects the complex model in favor of a U-Net (Solution 1) that preserves native spatial details and avoids relying on missing depth data, demonstrating skepticism towards superficially sophisticated but methodologically flawed approaches.
@@
To validate the model’s reasoning depth and provide transparency into our pipeline, we present four qualitative examples. First, we analyze two reasoning trajectories (Google Quest Challenge and TGS Salt Identification) to illustrate how the model acts as a skeptical critic, prioritizing methodological fit over superficial sophistication and overcoming human bias (Finding 5). Subsequently, we provide visual samples of two critical system artifacts: the Verbal Data Report and the Task Instruction, enabling a concrete inspection of the agent’s input and context.
@@ D.1 Case I: Overcoming Complexity Bias (Reasoning Analysis)
D.2 Case II: Domain Fit over Architectural Sophistication
To further substantiate the model’s robustness against complex-looking but flawed solutions (“Square peg in a round hole”), we present a second trajectory analysis in Figure 11.
Scenario and Conflict.
The agent evaluates two segmentation solutions for the TGS Salt Identification task:
  • Solution 0: A Vision Transformer (ViT-B/16) pre-trained on ImageNet.
  • Solution 1: A standard U-Net.
Human intuition might favor the ViT due to its top-tier status and global context modeling capabilities.
World Model Reasoning.
The World Model recognizes that the task requires pixel-perfect segmentation of small ($101\times 101$) seismic images. It correctly identifies that forcing inputs into the ViT’s required $224\times 224$ resolution introduces interpolation noise and destroys fine-grained salt details. Unswayed by the “SOTA” status, the model rejects the ViT in favor of the standard U-Net, citing the necessity of preserving native spatial resolution and avoiding severe domain mismatch.
D.3 Case III: Sample of the Verbal Data Report
@@
D.4 Case IV: Sample of the Task Instruction ($I$)